OpenStudy (anonymous):

Does anyone understand this question? Given: 2x – 3(x – 5 + x) = 2x – 3 Prove: x = 3 What are the missing justification steps, in order?

OpenStudy (anonymous):

Subtraction Property of Equality, Commutative Property of Addition, Addition Property of Equality Distributive Property, Commutative Property of Addition, Combine Like Terms Distributive Property, Associative Property of Addition, Combine Like Terms Subtraction Property of Equality, Associative Property of Addition, Addition Property of Equality

OpenStudy (anonymous):

2x – 3x +15 -3x - 2x + 3=0 -6x/6 = -18/6 x=3

OpenStudy (anonymous):

for the 1st one, there was simply a distribution. for the 2nd one, parens were added to "associate" two terms. for the 3rd one, like terms were combined

OpenStudy (anonymous):

So it's the last one?

OpenStudy (inkyvoyd):

@Answerthisnow

OpenStudy (inkyvoyd):

why don't you solve this equation step by step first?

OpenStudy (inkyvoyd):

Then, it will be much easier to find the justification...

OpenStudy (anonymous):

it is last one @Answerthisnow